# SceneComplete: Open-World 3D Scene Completion in Complex Real World Environments for Robot Manipulation

### v1

@@ Abstract
Careful robot manipulation in every-day cluttered environments requires an accurate understanding of the 3D scene, in order to grasp and place objects stably and reliably and to avoid mistakenly colliding with other objects. In general, we must construct such a 3D interpretation of a complex scene based on limited input, such as a single RGB-D image. We describe SceneComplete, a system for constructing a complete, segmented, 3D model of a scene from a single view. It provides a novel pipeline for composing general-purpose pretrained perception modules (vision-language, segmentation, image-inpainting, image-to-3D, and pose-estimation) to obtain high-accuracy results. We demonstrate its accuracy and effectiveness with respect to ground-truth models in a large benchmark dataset and show that its accurate whole-object reconstruction enables robust grasp proposal generation, including for a dexterous hand.
@@ I Introduction
As manipulation robots move from constrained environments such as factories to open-world environments such as homes and hospitals, they must be able to construct representations of their environment that enable robust, careful manipulation. Such representations need to individuate objects and characterize their shapes, so that the robot can reliably select stable grasps and placements for individual objects so that it can manipulate them without unwanted collisions. These representations must generally be constructed from limited input, such as a single RGB-D image. The problem is fundamentally ill-posed, but we are now in a position to address it using strong prior knowledge that has been learned by vision foundation models.
@@
In this paper, we propose a solution to this open-world scene completion problem in the form of a perception pipeline, SceneComplete. It combines multiple large pre-trained vision models into a system that takes a single RGB-D image as input and predicts as output a completed scene, consisting of a set of meshes for all the visible objects, including those that are partially occluded, Crucially, it makes no assumptions about the categories of the objects, their arrangement, or the camera viewpoint. It is constructed from multiple highly capable pre-trained perception components: a vision-language model (VLM) for identifying and generating short descriptions of the objects in a scene, a text-grounded image-segmentation model for localizing objects in the image, a 2D image-inpainting model for predicting the appearance of occluded parts of objects, an image-to-3D model for generating complete object meshes, and visual descriptor and pose-estimation modules to aid in composing individual predicted meshes into a final scene. None of these components can individually solve the problem, but in combination they provide robust object-centric interpretation of complex images, producing a segmented set of object meshes that are suitable for robot planning and manipulation.
As far as we know, this is the first system to support full-scene reconstruction from a single real-world RBG-D input in cluttered, occluded scenes with no assumptions about object categories. Many previous systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] have addressed simpler versions of this problem, with various combinations of known object categories, idealized sensory input, or lack of clutter. Handling the full generality of this problem class adds substantial difficulty to the problem.
@@
To demonstrate the effectiveness of SceneComplete, we conduct extensive quantitative and qualitative evaluations on real-world tabletop scenes. Our quantitative evaluations are on the Graspnet-1B dataset [12], which consists of heavily cluttered tabletop scenes. In these scenarios, accurately predicting the full scene—–including partially occluded objects—–is crucial for stable grasping, collision-free motion with an object in the hand, and reliable placing. We further illustrate the utility of our shape-reconstruction methods by using them as input to parallel-jaw [13] and dexterous grasping [14] proposal methods, the latter being especially sensitive to the detailed shape of the entire object.
II Method
Figure 2 illustrates the overall design of SceneComplete. It takes a single RGB-D image as input and produces a set of object meshes that are registered with the input 3D scan. The objective is to provide an accurate 3D reconstruction of the scene, in terms of segmentation into rigid components, and the shape of each component, expressed as a mesh. It operates in the following phases:
  1. 1.
    The RGB image is presented to a vision-language model, which is prompted to enumerate and describe the objects in the scene. (Sec. II-A)
  2. 2.
    These object descriptions, together with the RGB image, are input to a grounded segmentation model, which produces, for each object description, an image mask indicating the region the object occurs in. (Sec. II-A)
  3. 3.
    To handle occlusion, for each object, we mask out all other objects, and use an image-inpainting model to complete the 2D image of that object. (Sec. II-B)
  4. 4.
    Now, for each object, we have a 2D image corresponding to just that object, which we pass into an Image-to-3D model, which produces a high-quality textured mesh for the object, consistent with the image. (Sec. II-C)
  5. 5.
    The resulting object mesh has an arbitrary orientation, translation, and scale. For each object, we take the mesh and its associated image mask and partial point cloud (obtained in step 3) and determine an appropriate scale by finding correspondences and then solving for an appropriate scale. (Sec. II-D)
  6. 6.
    Finally, we determine an appropriate pose for each new mesh in the 3D coordinate frame of the original depth image using a 6DOF pose-estimation model. (Sec. II-E)
@@ II Method
Importantly, each of these steps makes use of existing pre-trained open-world visual-processing models, with almost no additional training (we have to do a small amount of low-rank adaptation of the inpainting model). This means that, as improved models become available for each of these tasks, as they inevitably will, we will be able to immediately profit from these improvements.
We describe each process in detail in the following subsections.
II-A Prompting and segmentation
@@
We begin by using a vision-language model to determine the number and basic description of objects in the scene. In our implementation, we pass the RGB image $I$ into ChatGPT-4o [15] with the prompt “describe the objects in the image with their generic name and color as prompts in a list. No description." It produces a text response from which we extract a list $t_{1},\ldots,t_{n}$ of text descriptions of objects. For example, in the example shown in Fig. 2, it returns “Blue Bowl, Tape, Banana" and so on.
@@
Next, we obtain an image mask for each object. For each text description $t_{i}$, we prompt a segmentation model (in our implementation, GroundedSam [16]) using $t_{i}$ on image $I$ and obtain a candidate set of pairs of masks and confidence values. For example, the prompt “a pear” might return multiple useful masks in a scene with two pears; in other cases, some of the masks will be unhelpful (and hopefully low-confidence). We use the confidence values to greedily select a set of non-overlapping masks, and associate each mask with the text prompt that generated it, producing the set $(R_{1},T_{1}),\ldots,(R_{N},T_{N})$, $R$ and $T$ denoting the masks and prompts, respectively, for $N$ objects in the scene. Our objective now is to compute a mesh corresponding to each of these hypothesized objects.
II-B Image Inpainting
@@
As illustrated in the Fig. 3, the objects represented by the masks in $(R_{1},T_{1}),\ldots,(R_{N},T_{N})$ could be partially occluded by other objects in the scene. In this step, we use an image inpainting algorithm to fill in the occluded parts of each object’s image. This is important for the next step of predicting a 3D mesh, because the image-to-3D models only perform reliably with a complete view of the object. Fig. 4 (a) illustrates the bad effects of attempting reconstruction from an incomplete image.
@@
In our implementation we begin with BrushNet [17], which takes an image with explicitly masked out regions and a text prompt, and produces a completed image, with the masked portions filled in. So for each $(R_{i},T_{i})$ pair, we begin by constructing an image $I_{i}$ with just the segmented object region $R_{i}$ on a white background. Then we construct an inpainting mask consisting of the union of the regions $R_{j}$ for $j\not=i$. Finally we query Brushnet with $I_{i}$, $\cup_{j\not=i}R_{j}$, and $T_{i}$ and obtain an inpainted image $P_{i}$ of the completed object on a white background.
@@
Figure 4: (a) The impact of inpainting on image-to-3D reconstruction. Without inpainting (top), the image-to-3D model generates incomplete meshes. Inpainting (bottom) fills in occluded parts, producing accurate 3D reconstructions. (b) Comparison of inpainting models. Unadapted BrushNet (middle) introduces artifacts, while the adapted version (right) correctly inpaints occluded objects producing a single fully observable object, improving 3D predictions for robotic manipulation in cluttered scenes.
@@
However, we observed that, out of the box, BrushNet occasionally synthesized additional objects in the occluded areas, as shown in Fig. 4 (b), possibly due to the “artistic” data set on which it was trained. To improve this behavior, we adapt BrushNet to full, single objects. It is important to note that, even though we want to adapt BrushNet to single fully observed objects, we want to retain the open world capabilities of the model. To achieve this, we use Low-Rank Adaptation (LoRA) [18] on its learnable layers with the PEFT [19] method, targeting domain-specific improvements on the tabletop YCB dataset [20]. LoRA is an effective method for adapting pretrained models to domain-specific outputs while retaining the inherent generalization of the models. To perform this adaptation, we project the different 3D YCB object meshes from arbitrary poses on a white background and add random brush masks for inpainting as suggested by BrushNet. The adapted model allows us to reliably inpaint individual occluded objects, including those from categories not in the adaptation dataset,
II-C Image-to-3D models for object reconstruction
@@
At this point, we have, for each object, a fully observed object image $P_{i}$ on a white background. The next step is to generate a 3D mesh model for the object. Although methods exist for operating directly on the point-cloud generated from the depth channel of the RGB-D image, the depth information is often very low-resolution and noisy, so these depth-only models tend to be highly tuned to specific categories [21, 22, 6, 2, 7] and viewpoints as shown in Fig. 7. There have been substantial recent improvements in RGB-only methods, which take advantage of the high quality of the RGB signal and enormous amounts of available training data [23, 24, 25, 26, 27, 28, 29].
@@
For RGB image as inputs, one option is to use methods that optimize a 3D mesh through differentiable rendering [30, 31, 32, 33, 34]. Although these models have impressive open world results, they have substantial computational cost. On the other hand, feed-forward methods such as InstantMesh [29] that directly map a single-view RGB image into a 3D mesh are highly accurate, open world, and computationally inexpensive at the time of inference. For these reasons, we use InstantMesh, providing each image $P_{i}$ as input and obtaining a complete textured 3D mesh $M_{i}$ as output. The resulting meshes are produced in an arbitrary orientation, at an arbitrary scale, so more work remains to be done, as explained in the subsequent subsections.
II-D Mesh Scaling using Dense Correspondence Matching
The next step is to rescale the meshes $M_{i}$, using a point-cloud constructed from the region $R_{i}$ to determine the scale as follows:
  • Using the default viewpoint $v$ from InstantMesh, generate an image $V_{i}$ as a 2D projection of mesh $M_{i}$.
  • Following [35], find dense visual descriptors of the segmented original image, $R_{i}$ and the projected image, $V_{i}$, using a pre-trained vision transformer (we use [36, 37]). In our observation, $v$ is usually close to the viewpoint from which $R_{i}$ was rendered. This enables us to obtain matching visual descriptors across $R_{i}$ and $V_{i}$.
  • Generate dense pixel-wise correspondences between the descriptor images, so that we have a set of pairs $(R_{i}^{j},V_{i}^{j})$ where $R_{i}^{j}$ is a pixel in the original image of the object and $V_{i}^{j}$ is a pixel in the synthesized view.
  • Map these correspondences into 3D, to obtain a pair $(R_{i}^{j^{\prime}},V_{i}^{j^{\prime}})$ of points in 3D. (Note that these point sets may not yet be aligned in 3D—we address that problem in the next step.)
  • Center each resulting 3D cloud, compute the average Euclidean distance from the points to the centroid in each one, and compute the ratio of these values as the scale factor.
  • Apply the scale factor to the mesh $M_{i}$ to put it in the same scale as the original point cloud to obtain $O_{i}$ as shown in Fig. 2-(5).
II-E 6D Object Pose Estimation and Registration
@@ II-D Mesh Scaling using Dense Correspondence Matching
As a result of this process, we have a set of pairs $(O_{i},\tau_{i})$, where each $O_{i}$ is a textured complete mesh for object $i$ and $\tau_{i}$ is a pose for that mesh in the camera coordinate frame. This scene reconstruction provides a highly general representation for a wide variety of downstream object-manipulation tasks. Accurate reconstructions of unobserved parts of objects enables a wide variety of manipulation operations, including many types of robust grasping, moving safely in cluttered but unobserved parts of the scene, moving safely when holding a grasped object, packing objects into boxes, etc.
III Experiments
Figure 5: Qualitative results on real-world RGB-D scene scans captured in our lab. For each scene we show: input RGB and depth, the completed scene from a different viewpoint, and a composite of the original point cloud with the completed scene, illustrating the accuracy and registration of the final result. These scenes, composed of arbitrary objects in various configurations, highlight the complexity of real-world environments. As can be seen, SceneComplete produces highly detailed 3D scene reconstructions in an open-world setting, despite significant occlusion and clutter, making it ideal for complex robotic manipulation tasks.
We evaluate the accuracy and utility of SceneComplete in three different regimes:
  • Quantitative evaluation of reconstruction accuracy on a large-scale dataset of tabletop scenes with RGB-D scans captured using a Realsense D435 camera, GraspNet-1B [12], against the ground-truth object models and poses.
  • Qualitative evaluation of reconstruction on a smaller-scale dataset of tabletop scenes with RGB-D images (also collected with a Realsense D435) but using objects in our lab that have little chance of having appeared in the training of any of the models in the system.
  • Task-driven evaluation of reconstruction by demonstrating that the reconstructed object models are sufficiently accurate to enable grasp proposals for a multi-fingered hand, which depends on having a good estimate of the entire object shape.

### v4

# SceneComplete: Open-World 3D Scene Completion in Cluttered Real World Environments for Robot Manipulation

@@ Abstract
Careful robot manipulation in every-day cluttered environments requires an accurate understanding of the 3D scene, in order to grasp and place objects stably and reliably and to avoid colliding with other objects. In general, we must construct such a 3D interpretation of a complex scene based on limited input, such as a single RGB-D image. We describe SceneComplete, a system for constructing a complete, segmented, 3D model of a scene from a single view. SceneComplete is a novel pipeline for composing general-purpose pretrained perception modules (vision-language, segmentation, image-inpainting, image-to-3D, visual-descriptors and pose-estimation) to obtain highly accurate results. We demonstrate its accuracy and effectiveness with respect to ground-truth models in a large benchmark dataset and show that its accurate whole-object reconstruction enables robust grasp proposal generation, including for a dexterous hand. We release the code and additional results on our website.
@@ I INTRODUCTION
As manipulation robots move from constrained environments such as factories and workshops to open-world environments such as homes and hospitals, they must be able to construct representations of their environment that enable robust, careful manipulation. Such representations need to individuate objects and characterize their shapes, so that the robot can reliably select stable grasps and placements for individual objects and manipulate them without unwanted collisions. These representations must generally be constructed from limited input, such as a single RGB-D image. This problem is fundamentally ill-posed, but we are now in a position to address it using strong priors that have been learned by vision foundation models.
@@
In this paper, we propose a solution to this open-world scene completion problem in the form of a perception pipeline, SceneComplete. It combines multiple large pre-trained vision models into a system that takes a single RGB-D image as input and predicts as output a completed scene, consisting of a set of meshes for all the visible objects, including those that are partially occluded. Crucially, it makes no assumptions about the categories of the objects, their arrangement, or the camera viewpoint. It is constructed from multiple highly capable pre-trained perception components: a vision-language model (VLM) for identifying and generating short descriptions of the objects in a scene, a text-grounded image-segmentation model for localizing objects in the image, a 2D image-inpainting model for predicting the appearance of occluded parts of objects, an image-to-3D model for generating complete object meshes, and visual descriptor and pose-estimation modules to aid in composing individual predicted meshes into a final scene. None of these components can individually solve the problem, but in combination they provide robust object-centric interpretation of complex images, producing a segmented set of object meshes that are suitable for robot planning and manipulation.
To demonstrate the effectiveness of SceneComplete, we conduct extensive quantitative and qualitative evaluations on real-world tabletop scenes. Our quantitative evaluations are on the Graspnet-1B [1] and YCB-Video [2] datasets, which consist of cluttered tabletop scenes. In these scenarios, accurately predicting the full scene—–including partially occluded objects—–is crucial for stable grasping, collision-free motion with an object in the hand, and reliable placing.
We further illustrate the utility of our shape-reconstruction methods by using them as input to parallel-jaw [3, 4] and dexterous grasping [5] methods, the latter being especially sensitive to the detailed shape of the entire object.
II Related Work
Feed-forward Scene Reconstruction: Feed-forward multi-object scene completion methods such as  [6, 7, 8, 9, 10] learn end-to-end mappings from single-view RGB-D input to completed object meshes or occupancy grids. These methods are attractive because they are computationally fast at inference time; however, suffer from key limitations: ShAPO [7], FSD [6], and CRISP [10] are closed-set methods, trained extensively on fixed benchmark datasets with predefined object categories, and fail to generalize to novel objects. OctMAE [8] addresses open-set scenes by performing scene-level reconstruction. However, its predictions are surface-level and do not individuate objects, limiting their utility for robotic manipulation. ZeroGrasp [9] performs scene reconstruction with individuated objects for grasp generation, but the reconstruction quality—crucial for collision avoidance—is subpar since the focus is primarily on grasping. In our results, we compare and outperform OctMAE and ZeroGrasp in reconstructing scenes and generating grasps.
Compositional Scene Reconstruction: This line of work composes multiple open-set models with zero-shot generalization into a pipeline for scene reconstruction. While such methods are typically slower than feedforward methods, they can directly handle unseen objects with little or no retraining [11, 12, 13]. CAST [11] and Gen3DSR [12] both reconstruct full scenes from a single RGB image by first predicting depth maps from monocular depth estimation models. CAST then generates meshes for individual objects and applies a physics-aware correction step to enforce physically consistent placements, while Gen3DSR integrates components such as DreamGaussian to produce full 3D meshes. However, because both methods rely on predicted depth, their pipelines are tightly coupled to synthetic outputs and fail to adapt to raw RGB-D input, where sensor noise is unavoidable. These approaches are primarily designed for asset generation, and their inability to incorporate ground-truth depth limits their applicability in robotics. By contrast, SceneComplete leverages observed sensor depth (often noisy), enabling a more flexible pipeline tailored for manipulation tasks.
A different line of work, Open6DOR [13], introduces a benchmark for language-driven 6-DoF object rearrangement, along with a baseline pipeline for grasp generation. Their approach composes a set of modules similar to those in SceneComplete, but focuses exclusively on predicting object poses. As a result, the pipeline does not perform full scene reconstruction, omits critical steps such as registration, and restricts predictions to objects that are almost fully visible.
@@ II Related Work
III Method
@@
Figure 2 illustrates the overall design of SceneComplete. It takes a single RGB-D image as input and produces a set of object meshes that are registered with the input 3D scan. The objective is to provide an accurate 3D reconstruction of the scene, in terms of segmentation into rigid components, and the shape of each component, expressed as a mesh. Importantly, each step in the pipeline makes use of existing pre-trained open-world visual-processing models, with almost no additional training (we do a small amount of low-rank adaptation of the inpainting model). This means that, as improved models become available for each of these tasks, as they inevitably will, we will be able to immediately profit from these improvements. We describe each process in detail in the following subsections.
III-A Prompting and segmentation
@@
We begin by using a vision-language model to determine the number and basic description of objects in the scene. In our implementation, we pass the RGB image $I$ into ChatGPT-4o [^1] https://openai.com/index/hello-gpt-4o/ with the prompt “describe the objects in the image with their generic name and color as prompts in a list.” It produces a text response as a list $t_{1},\ldots,t_{n}$ of text descriptions of objects. E.g., in the example shown in Fig. 2, it returns “Blue Bowl, Tape, Banana” and so on.
@@
Next, we obtain an image mask for each object. For each text description $t_{i}$, we prompt a grounded segmentation model (in our implementation, GroundedSam2 [14]) using $t_{i}$ on image $I$ and obtain a candidate set of pairs of masks and confidence values. For example, the prompt “a pear” might return multiple useful masks in a scene with two pears; in other cases, some of the masks will be unhelpful (and hopefully low-confidence). We use the confidence values to greedily select a set of non-overlapping masks, and associate each mask with the text prompt that generated it, producing the set $(R_{1},T_{1}),\ldots,(R_{N},T_{N})$, $R$ and $T$ denoting the masks and prompts, respectively, for $N$ objects in the scene.
III-B Image Inpainting
@@
As illustrated in the Fig. 3, the objects represented by the masks in $(R_{1},T_{1}),\ldots,(R_{N},T_{N})$ could be partially occluded by other objects in the scene. In this step, we use an image inpainting algorithm to fill in the occluded parts of each object’s image. This is important for the next step of predicting a 3D mesh, because the image-to-3D models only performs reliably with a complete view of the object. Fig. 4 (a) illustrates the poor results of attempting reconstruction from an incomplete image.
@@
In our implementation we begin with BrushNet [15], which takes an image with explicitly masked out regions and a text prompt, and produces a completed image, with the masked portions filled in. So for each $(R_{i},T_{i})$ pair, we begin by constructing an image $I_{i}$ with just the segmented object region $R_{i}$ on a white background. Then we construct an inpainting mask consisting of the union of the regions $R_{j}$ for $j\not=i$. Finally we query Brushnet with $I_{i}$, $\cup_{j\not=i}R_{j}$, and $T_{i}$ and obtain an inpainted image $P_{i}$ of the completed object.
@@
Figure 4: (a) The impact of inpainting on image-to-3D reconstruction. Without inpainting (top), the image-to-3D model generates incomplete meshes. Inpainting (bottom) fills in occluded parts, producing accurate 3D reconstructions. (b) Comparison of inpainting models. Unadapted BrushNet (middle) introduces artifacts, while the adapted version (right) inpaints occluded parts correctly producing a fully observed object.
@@
However, we observed that, out of the box, BrushNet occasionally synthesized additional objects in the occluded areas, as shown in Fig. 4 (b), possibly due to the “artistic” data set on which it was trained. To improve this behavior, we adapt BrushNet to full, single objects. It is important to note that, even though we want to adapt BrushNet to single fully observed objects, we want to retain the open world capabilities of the model. To achieve this, we use Low-Rank Adaptation (LoRA) [16] on its learnable layers with the PEFT method, targeting domain-specific improvements on the tabletop YCB dataset [17]. LoRA is an effective method for adapting pretrained models to domain-specific outputs while retaining the inherent generalization of the models. To perform this adaptation, we project the different 3D YCB object meshes from arbitrary poses on a white background and add random brush masks for inpainting as suggested by BrushNet. The adapted model allows us to reliably inpaint individual occluded objects, including those from categories not in the adaptation dataset.
III-C Image-to-3D models for object reconstruction
@@
At this point, we have, for each object, a fully observed object image $P_{i}$ on a white background. The next step is to generate a 3D mesh model for the object. Although methods exist for operating directly on the point-cloud generated from the depth channel of the RGB-D image, the depth information is often very low-resolution and noisy, so these depth-only models tend to be highly tuned to specific categories [18, 19, 7, 20] and viewpoints, or operate over idealized sensory input. In recent years, substantial improvements have been made in RGB-only methods, which take advantage of the high quality of the RGB signal and the enormous amounts of available training data [21, 22, 23].
@@
For RGB image as inputs, one option is to use methods that optimize a 3D mesh through differentiable rendering [24, 25, 26]. Although these models have impressive open world results, they have substantial computational cost. On the other hand, feed-forward methods such as InstantMesh [23] that directly map a single-view RGB image into a 3D mesh are highly accurate, open world, and computationally inexpensive at the time of inference. For these reasons, we use InstantMesh, providing each image $P_{i}$ as input and obtaining a complete textured 3D mesh $M_{i}$ as output. The resulting meshes are produced in an arbitrary orientation, at an arbitrary scale, so more work remains to be done, as explained in the subsequent subsections.
III-D Mesh Scaling using Dense Correspondence Matching
The next step is to rescale the meshes $M_{i}$, using a point-cloud constructed from the region $R_{i}$ to determine the scale factor as follows: (i) Using the default viewpoint $v$ from InstantMesh, generate an image $V_{i}$ as a 2D projection of mesh $M_{i}$. (ii) Following [27], find dense visual descriptors of the segmented original image, $R_{i}$ and the projected image, $V_{i}$, using a pre-trained vision transformer. In our observation, $v$ is usually close to the viewpoint from which $R_{i}$ was rendered. This enables us to obtain matching visual descriptors across $R_{i}$ and $V_{i}$. (iii) Generate dense pixel-wise correspondences between the descriptor images, so that we have a set of pairs $(R_{i}^{j},V_{i}^{j})$ where $R_{i}^{j}$ is a pixel in the original image of the object and $V_{i}^{j}$ is a pixel in the synthesized view. (iv) Map these correspondences into 3D, to obtain a pair $(R_{i}^{j^{\prime}},V_{i}^{j^{\prime}})$ of points in 3D. (Note that these point sets may not yet be aligned in 3D—we address that problem in the next step.) (v) Center each resulting 3D cloud, compute the average Euclidean distance from the points to the centroid, and compute the ratio of these values as the scale factor. (vi) Apply the scale factor to the mesh $M_{i}$ to put it in the same scale as the original point cloud to obtain $O_{i}$ as shown in Fig. 2-(5).
III-E 6D Object Pose Estimation and Registration
@@ III-D Mesh Scaling using Dense Correspondence Matching
As a result of this process, we have a set of pairs $(O_{i},\tau_{i})$, where each $O_{i}$ is a textured complete mesh for object $i$ and $\tau_{i}$ is a pose for that mesh in the camera coordinate frame. This scene reconstruction provides a highly general representation for a wide variety of downstream object-manipulation tasks. Accurate reconstructions of unobserved parts of objects enables a wide variety of manipulation operations, including many types of robust grasping, moving safely in cluttered but unobserved parts of the scene, moving safely when holding a grasped object, etc.
IV Experiments
Figure 5: Qualitative comparisons of scene reconstructions on the GraspNet-1B dataset. For each scene we show, the input RGB-D image, OctMAE reconstruction (rendered as normal maps as it predicts scene-level occupancy values), ZeroGrasp reconstruction (rendered as normal maps), our reconstruction (visualized as individually reconstructed object meshes color-matched to the ground truth), and ground-truth object meshes. Highlighted regions indicate missing area (black) or spurious region connecting distinct objects (red).
Figure 6: Evaluating GSR on the YCB-V dataset in Isaac Gym. We show a close-up of 3 distinct objects being picked up.
Figure 7: We demonstrate dexterous grasps using both Shadow Hands and Allegro Hands [5] on objects from the GraspNet-1B dataset, highlighting improved manipulation of complex objects with complete 3D reconstructions.
We evaluate SceneComplete through three main experimental regimes:
  • Scene Reconstruction and Grasping: We compare SceneComplete against an existing single-view scene reconstruction method on a large-scale dataset of tabletop scenes, GraspNet-1B [1], captured using a RealSense D435 camera. We also evaluate how the reconstructed scenes contribute in generating collision-free grasps.
  • Object-grasping and Dexterous Manipulation: We evaluate the effectiveness of SceneComplete in generating grasps for successfully picking up objects inside a simulation environment using a parallel-jaw gripper. We also demonstrate that the reconstructed object models are of sufficient fidelity to enable dexterous grasp proposals for a multi-fingered hand, which depends on having a good estimate of the entire object shape.
  • Real-world Evaluations: To assess the real-world applicability of SceneComplete, we conduct pick-and-place experiments on a physical robot on a smaller-scale dataset of tabletop scenes collected in our lab (also using a RealSense D435 camera), that include everyday objects which are less likely to have appeared in the training distribution of any of the models used. We evaluate across 15 scenes, each with 4 to 6 objects.